OpenStudy (dtan5457):

Calculate the standard heat of formation of Cu2O(s) given the reactions 1. 2Cu2O(s)+O2(g)>>>4CuO(s) delta H<<-288.0kj 2. CuO(s)+Cu(s)>>>Cu2O(s) delta H<<<-11.3 kJ

OpenStudy (dtan5457):

@Abhisar @aaronq

OpenStudy (aaronq):

The Hf's for O2 and Cu are zero, so you have: 1) \( \sf ΔH_1=4H_f(CuO)−2H_f(Cu2O)\) 2) \(\sf ΔH_2=H_f(Cu2O)−H_f(CuO)\) Now it's a matter of solving the system of equations 1) \(\sf \dfrac{ΔH_1+2H_f(Cu_2O)}{4}=H_f(CuO)\) Plug into 2) 2) \( \sf ΔH_2=H_f(Cu_2O)−\dfrac{ΔH_1+2H_f(Cu_2O)}{4}\) Rearrange to solve for \(H_f(Cu_2O)\) 2) \( \sf ΔH_2+\dfrac{ΔH_1}{4}=H_f(Cu_2O)-2H_f(Cu_2O)\) \(\sf H_f(Cu_2O)=-[ΔH_2+\dfrac{ΔH_1}{4}]\)

OpenStudy (dtan5457):

Forgot about this question..got it now. thanks